OpenStudy (anonymous):

a triangle has three angles. the second angle is 20 degrees more than the first. the third angle is 8 degrees more than the second. find all three angles.

OpenStudy (anonymous):

What do the angles of a triangle add up to?

OpenStudy (anonymous):

that's all the information it gave me.

OpenStudy (anonymous):

The interior angles of a triangle always add up to 180. We know that one angle is 20 degrees more than the first angle, and the third is 8 degrees more than the second. If we say that angle 1 has a value of "a," we can write" Angle 1 = a Angle 2 = a + 20 Angle 3 = a + 20 + 8 = a +28 Angle 1 + Angle 2 + Angle 3 = 180 a + a + 20 + a +28 = 180 Solve for a, and you have the first angle. Plug that value in for a for the other two, and you'll know all three.

OpenStudy (anonymous):

why do you have 3 "a"s @Vandreigan

OpenStudy (anonymous):

a is the value of the first angle. The second angle is 20 degrees more than this, so it's a+20. The third is 8 more than the second, which is a+28. If you add up all three of the angles for a triangle, you get 180 (always). The last equation, the one with the three a's, is me adding up all the angles (a, a+20, and a+28) and setting it equal to 180 so that we can find a (the first angle), which will also tell us the other two angles.

OpenStudy (anonymous):

a is 44, thanks for you help!! @Vandreigan

OpenStudy (anonymous):

There ya go! :) My pleasure